# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning

### v1

@@ V Experiments
    Swin-Base (U) means using the Swin-Base backbone plus UperNet segmentation algorithm. RingMo-Aerial (U), (A), (M) represent the RingMo aerial backbone with UperNet segmentation algorithm, ARS-Adapter and UperNet algorithm, and Mask2Former algorithm. Same below.
TABLE V: Comp. on UDD6
@@
TABLE VI: Comp. on Floodnet
@@ V-A Pre-train
In the pretraining phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized. Prior to pretraining, we initialize the model parameters using Kaiming initialization. Subsequently, we perform 100 epochs of masked image modeling (MIM) pretraining. During MIM pretraining, the input image size is set to 224, and random cropping and scaling are applied to the input images. The mask size is set to 32 pixels. The batch size is set to 2048, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 2 epochs. Following MIM pretraining, we continue with contrastive learning (CL) pretraining, jointly training for 36 epochs. The input image size remains 224, and similar random cropping and scaling are applied. The batch size is adjusted to 1024, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 1 epoch.
V-B Scene Classification
V-B1 Datasets and Evaluation Metrics
For semantic segmentation, the AID [94] and RESISC-45 [95] datasets are employed.
  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 different types of scenes.
In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively. In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively. The experiments use overall accuracy (OA) as a performance metric, which is the most commonly used performance metric in scene classification.
V-B2 Experiment Settings
The experiment of the scene classification is based on mmpretrain , Cascade RCNN [96] is employed as the classification framework. The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
V-B3 Comparison With the SOTA
The comparison results are shown in Table II. RingMo-Aerial combined with Cascade RCNN can achieve classification results far exceeding other methods.
V-C Change Detection
V-C1 Datasets and Evaluation Metrics
  • LEVIR. The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset. It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5 meter resolution. Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels. Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases, respectively.
  • CDD. The CDD dataset is a comprehensive collection of bitemporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels. These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches. This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs. To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
V-C2 Experiment Settings
Our model adopts BIT with Swin Transformer as the backbone. The experiments use overall accuracy (OA) and F1 and IoU scores with regard to the change detection as the primary evaluation metrics. The model is trained with AdamW optimizer with a weight decay 0.01 and beta values equal to (0.9, 0.999). The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size of 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
V-C3 Comparison With the SOTA
Figure 12: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
The comparison results are shown in Table III. RingMo-Aerial combined with BIT can achieve change detection results far exceeding other methods. In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection. When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated in our architecture.
The comparison of visual results12 on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks. The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images. In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
RingMo-Aerial’s commendable performance not only makes it a key player in the field, but also highlights the importance of the impact of foundation model innovation. These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
V-D 3D-Reconstruction
V-D1 Datasets and Evaluation Metrics
For 3D-Reconstruction, the and WHU [99] datasets are employed.
V-D2 Experiment Settings
Our work is coded based on the PyTorch framework. The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1. During the training process, our method uses the Adam optimizer[101] for parameter optimization, and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001. In addition, our method is consistent with the existing multi-stage networks[102]. First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively. Then, deep optimization is performed through three stages of deep hypotheses, and the number of hypotheses is 48, 32, and 8, respectively. During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
V-D3 Comparison With the SOTA
TABLE VII: Performance comparisons on the LuoJia-MVS and WHU datasets.
Number of Views Model LuoJia-MVS WHU
MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
Three views * PatchmatchNet[103] 0.252 92.7 87.2 0.173 96.5 94.8
Fast-MVSNet[104] 0.194 95.7 92.0 0.184 95.5 94.1
MVSNet[105] 0.172 96.1 92.4 0.190 95.0 94.3
R-MVSNet[106] 0.177 96.0 93.5 0.183 95.3 93.5
RED-Net[99] 0.109 98.2 96.9 0.112 98.1 97.9
Cas-MVSNet[102] 0.103 98.4 97.1 0.111 97.7 97.6
RingMo-Aerial 0.095 98.7 97.8 0.121 97.8 97.5
Five views * PatchmatchNet[103] 0.283 90.4 84.1 0.160 96.9 95.0
Fast-MVSNet[104] 0.357 84.6 74.9 0.157 96.1 95.6
MVSNet[105] 0.270 91.2 81.8 0.160 95.8 95.5
R-MVSNet[106] 0.259 92.3 86.7 0.173 95.4 93.8
RED-Net[99] 0.156 94.9 90.5 0.104 98.1 97.9
Cas-MVSNet[102] 0.141 97.9 95.4 0.095 97.8 97.8
RingMo-Aerial 0.122 98.1 96.5 0.091 98.0 97.9
    ‘Three Views’ and ‘Five Views’ represent the number of views for the input of the network is three and five, respectively.
Table VII provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets. Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error. On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet, and improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs. 0.141), 0.2% (98.1% vs. 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a large improvement. In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%. It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
V-E Semantic Segmentation
V-E1 Datasets and Evaluation Metrics
  • UAVID. The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid and 15 test), and each sequence contains 10 pictures. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
  • UDD6. The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images. All of the images are 4K resolution. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. The training set is used for training and the validation set for testing.
In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc) and average accuracy (aAcc).
V-E2 Experiment Settings
The experiment of the semantic segmentation is based on mmsegmentation [110], UperNet [111] and Mask2Former [112] are employed as the segmentation framework. The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
V-E3 Comparison With the SOTA
The comparison results on the UAVID dataset are shown in Table IV, results on the UDD6 dataset are shown in Table V, results on the FloodNet dataset are shown in Table VI. RingMo-Aerial combined with Mask2Former can achieve segmentation results far exceeding other methods. Experiments on UperNet (RingMo-Aerial vs. Swin-Base) also show that RingMo-Aerial is a more suitable basic model for ARS segmentation tasks.
@@ V-E3 Comparison With the SOTA
Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks. As illustrated in the Figure11, RingMo-Aerial consistently produces superior segmentation maps. Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures. In particular, the RingMo-Aerial + UperNet + ARS-Adapter and RingMo-Aerial + Mask2Former methods demonstrate higher accuracy and clarity in the segmentation of small targets (e.g., stationary and moving vehicles), with enhanced boundary and detail capture. This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
V-F Object Detection
TABLE VIII: Comp. on VisDrone-DET
@@ V-F Object Detection
TABLE IX: Comp. on UAVDT
@@
TABLE X: Comp. on ShipDataset
@@
TABLE XI: Comp. on DroneVehicle
@@
Figure 13: Qualitative results for the VisDrone-DET dataset when using different methods.
TABLE XII: Comp. on SARData
@@
RingMo-Aerial(Y) * RingMo-Aerial 34.0 81.5
RingMo-Aerial(C) * RingMo-Aerial 34.9 82.4
@@
    RingMo-Aerial (A), (Y), (C) represent the RingMo aerial backbone with ARS-Adapter and YOLOX-S algorithm, YOLOX-S algorithm, and Cascade RCNN algorithm.
V-F1 Datasets and Evaluation Metrics
  • UAVDT. The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with the average resolution of 1080 × 540 pixels. The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
  • SARData. To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside. This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection. The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels. The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
V-F2 Experiment Settings
The experiment of the object detection is based on mmdetection [148], Cascade RCNN [96], YOLOX-S[134] and DDQ DETR [127] are employed as the detection framework. The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used). That for YOLOX-S spanned 36 epochs with a batch size of 4 total (4 per GPU and 1 NVIDIA A40 GPU are used). That for DDQ DETR spanned 36 epochs with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the learning rate is decreased by a factor of 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
TABLE XIII: Object Tracking: Comp. on Visdrone-MOT

### v3

@@ 4.4 Downstream Fine-tuning
    Swin-Base (U) uses the Swin-Base backbone plus the UperNet segmentation algorithm. RingMo-Aerial (U), (A), (M) represent the RingMo aerial backbone with UperNet segmentation algorithm, ARS-Adapter and UperNet algorithm, and Mask2Former algorithm. Same below.
In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks. Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model. The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, shown as Fig.12. The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking. Like AdaptFormer [55], the introduced ARS-Adapter module operates in parallel with the model. The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer. Therefore, the forward propagation is:
[…]
Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling. In parallel with FFN, the adapter introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
[…]
5 Experiments
TABLE II: Semantic Segmentation: Comp. on UDD6
@@ 5 Experiments
TABLE III: Semantic Segmentation: Comp. on Floodnet
@@ 5.1 Pre-training Implementation
In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized. Before pre-training, we initialize the model parameters using Kaiming initialization. Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training. During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images. The mask size is set to 32 pixels. The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs. Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs. The input image size remains 224, and similar random cropping and scaling are applied. The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
5.2 Performance on Single/Multi-modal Downstream Tasks
We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios. Detailed descriptions of the datasets, experimental setups, and additional visualizations for each task are provided in the Appendix.
5.2.1 Semantic Segmantation
Comparative results on the UAVID, UDD6, and FloodNet datasets are presented in Tabs. I, II, and III, respectively. The combination of RingMo-Aerial with Mask2Former achieves segmentation outcomes that significantly outperform those of other methods.
@@ 5.2.1 Semantic Segmantation
Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks. RingMo-Aerial consistently produces superior segmentation maps. Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures. In particular, the RingMo-Aerial + UperNet + ARS-Adapter and RingMo-Aerial + Mask2Former methods demonstrate higher accuracy and clarity in the segmentation of small targets (e.g., stationary and moving vehicles), with enhanced boundary and detail capture. This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
5.2.2 Object Detection
TABLE IV: Object Detection: Comp. on VisDrone-DET
@@ 5.2.2 Object Detection
TABLE V: Object Detection: Comp. on UAVDT
@@
TABLE VI: Object Detection: Comp. on ShipDataset
@@
TABLE VII: Object Detection: Comp. on DroneVehicle
@@
TABLE VIII: Object Detection: Comp. on AerialSARData
@@
RingMo-Aerial(C) * RingMo-Aerial 34.9 82.6
@@
    RingMo-Aerial (A), (C) represent the RingMo-Aerial backbone with ARS-Adapter and Cascade RCNN algorithm.
An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance. On unimodal datasets such as VisDrone-DET (as shown in Tab.IV) and UAVDT (as shown in Tab.V), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods. The experimental results further demonstrate that RingMo-Aerial, compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.VI), and PP-PicoDet-L follows. In contrast, YOLOX-L and YOLOX-Tiny show the worst results. Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
For the multimodal dataset, DroneVehicle (as shown in Tab.VII), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods. Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial. This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications. Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.VIII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance. Mainly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and a mAP50 of 84.4%, highlighting its superiority in SAR modality. These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery. Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments. These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields. For optical images, as shown in the appendix, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets. In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups. Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms. RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images. As shown in the appendix, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence. The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images. RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for ARS.
5.2.3 Scene Classification
TABLE IX: Scene Classification: Comp. on AID and RESISC-45
Model AID (OA) RESISC-45 (OA)
(TR=20%/50%) (TR=10%/20%)
GASSL [96] 93.55/95.92 90.86/93.06
SeCo [33] 93.47/95.99 89.64/92.91
SatMAE [97] 95.02/96.94 91.72/94.10
RingMo [1] 96.90/98.34 94.25/95.67
RVSA [98] 97.03/98.50 93.93/95.69
DINO-MC [99] 95.16/97.09 -
TOV [100] 95.16/97.09 90.97/93.79
SSL4EO [101] 91.06/94.74 87.60/91.27
CMID [102] 96.11/97.79 91.84/94.07
CACo [103] 90.88/95.05 88.28/91.94
CROMA [104] 96.44/97.58 92.63/95.04
SatLas [105] 94.96/97.38 92.16/94.70
GFM [106] 95.47/97.09 -
Scale-MAE [107] 96.44/97.58 92.63/95.04
SkySense [2] 97.68/98.60 94.85/96.32
RingMo [1] 96.90/98.34 94.25/95.67
RingMo-Aerial 95.81/96.46 92.28/95.65
The comparison results are shown in Tab.IX. Although RingMo-Aerial does not perform as well as specialized satellite models like SkySense in nadir-view classification, it still demonstrates strong performance, reflecting its robust generalization capability.
5.2.4 Change Detection
TABLE X: Change Detection: Comp. on LEVIR-CD and CDD
Methods LEVIR-CD CDD
F1 IoU F1 IoU
FC-Siam-Di [108] 86.31 83.31 70.61 54.57
FC-Siam-Conc [108] 83.69 76.77 75.11 60.14
IFNet [109] 88.13 82.93 90.30 71.91
STANet [110] 87.30 77.40 84.12 72.22
BIT [111] 89.31 89.37 88.90 80.01
SNUNet [112] 88.16 87.17 83.89 72.11
ChangeFormer [113] 90.40 88.80 89.83 81.53
RingMo [1] 89.53 82.38 90.90 84.09
RingMo-Aerial 92.71 87.10 94.62 90.09
The comparison results are shown in Tab.X. RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods. In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection. When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in the appendix. The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images. In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
RingMo-Aerial’s commendable performance not only makes it a key player in the field but also highlights the importance of the impact of foundation model innovation. These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
5.2.5 3D-Reconstruction
TABLE XI: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
Number of Views Model LuoJia-MVS WHU
MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
Three views * PatchmatchNet[114] 0.252 92.7 87.2 0.173 96.5 94.8
Fast-MVSNet[115] 0.194 95.7 92.0 0.184 95.5 94.1
MVSNet[116] 0.172 96.1 92.4 0.190 95.0 94.3
R-MVSNet[117] 0.177 96.0 93.5 0.183 95.3 93.5
RED-Net[118] 0.109 98.2 96.9 0.112 98.1 97.9
Cas-MVSNet[119] 0.103 98.4 97.1 0.111 97.7 97.6
RingMo-Aerial 0.095 98.7 97.8 0.121 97.8 97.5
Five views * PatchmatchNet[114] 0.283 90.4 84.1 0.160 96.9 95.0
Fast-MVSNet[115] 0.357 84.6 74.9 0.157 96.1 95.6
MVSNet[116] 0.270 91.2 81.8 0.160 95.8 95.5
R-MVSNet[117] 0.259 92.3 86.7 0.173 95.4 93.8
RED-Net[118] 0.156 94.9 90.5 0.104 98.1 97.9
Cas-MVSNet[119] 0.141 97.9 95.4 0.095 97.8 97.8
RingMo-Aerial 0.122 98.1 96.5 0.091 98.0 97.9
    ‘Three Views’ and ‘Five Views’ represent the number of views for the network’s input is three and five, respectively.
Tab.XI provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets. Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error. On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet. It improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs. 0.141), 0.2% (98.1% vs. 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a significant improvement. In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%. It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
5.2.6 Object Tracking
TABLE XII: Object Tracking: Comp. on Visdrone-MOT